OpenStudy (anonymous):

Find the least common denominator for these two rational expressions 3/n^2 -2/5n

OpenStudy (anonymous):

Is it written \[\frac{ 3 }{ n }^2\] or \[\frac{ 3 }{ n^2 }\]?

OpenStudy (anonymous):

@Brostep0s the second one

OpenStudy (anonymous):

\[\frac{ 2 }{ 5n^2 }\]?

OpenStudy (anonymous):

I found this answer: 5n^2. Is this correct?